OpenStudy (anonymous):

For what value of θ, in degrees, is sin θ = cos 58°

OpenStudy (mathstudent55):

Look at the figure below. 

OpenStudy (mathstudent55):

Using \(\sin \theta = \dfrac{opp}{hyp}\), can you tell what is sin A?

Parth (parthkohli):

\[\cos(90^{\circ}- \theta )= \cos (58^{\circ}) \]\[\Rightarrow 90^\circ - \theta = 360^\circ \pm 58^\circ\]